### v1

# Gas-phase metallicity gradients in galaxies at Key Words.: Galaxies: high-redshift ā Galaxies: abundances ā Galaxies: ISM ā Galaxies: evolution ā Techniques: imaging spectroscopy ā Techniques: high angular resolution1 Introduction

The study of gas-phase metallicity and its spatial distribution at high redshift is crucial to understand the processes that shaped the growth and evolution of galaxies in the early Universe. Here we study the spatially resolved metallicity in three systems at 6ā8, namely A2744-YD4, BDF-3299, and COSMOS24108, with JWST NIRSpec IFU low-resolution ( 100) spectroscopic observations. These are among the highest- sources in which metallicity gradients have been probed so far. Each of these systems hosts several spatial components in the process of merging within a few kiloparsecs, identified from the rest-frame UV and optical stellar continuum and ionised gas emission line maps. The sources have heterogeneous properties, with stellar masses log() 7.6ā9.3, star formation rates (SFRs) 1ā15 Ā yr, and gas-phase metallicities 12+log(O/H) 7.7ā8.3, which exhibit a large scatter within each system. Their properties are generally consistent with those of the highest-redshift samples to date ( 3ā10), though the sources in A2744-YD4 and COSMOS24108 are at the high end of the mass-metallicity relation (MZR) defined by the 3ā10 sources. Moreover, the targets in this work follow the predicted slope of the MZR at 6ā8 from most cosmological simulations. The gas-phase metallicity gradients are consistent with being flat in the main sources of each system. Flat metallicity gradients are thought to arise from gas mixing processes on galaxy scales, such as mergers or galactic outflows and SN winds driven by intense stellar feedback, which wash out any gradient formed in the galaxy. The existence of flat gradients at 6ā8 sets also important constraints on cosmological simulations, whose predictions on the cosmic evolution of metallicity gradients often differ significantly, especially at high redshift.
Gas-phase metallicity gradients in galaxies at
Key Words.:
Galaxies: high-redshift ā Galaxies: abundances ā Galaxies: ISM ā Galaxies: evolution ā Techniques: imaging spectroscopy ā Techniques: high angular resolution1 Introduction
Gas-phase metallicity, its scaling relations with other galactic properties, and its spatial distribution are fundamental tools to study and understand the evolution of galaxies (e.g. Maiolino & Mannucci, 2019). The interplay between gas accretion, star formation, outflows, and mergers regulates the growth of galaxies and the build up of their metal content, leading to the correlation between gas-phase metallicity and stellar mass in galaxies known as mass-metallicity relation (MZR; e.g. Davé et al. 2011; Pallottini et al. 2014; Somerville & Davé 2015). The spatial distribution of metals in galaxies, usually described through radial metallicity gradients, bears the imprint of these underlying processes. Therefore, spatially resolved studies of metallicity in the early Universe, where the formation and primordial growth of galaxies were taking place, are key to better constrain the main mechanisms that have contributed to shape them.
Negative (radially decreasing) gradients are usually interpreted as resulting from an inside-out galaxy formation scenario, in which stars start forming earlier in the inner parts of galaxies and thus have more time to chemically enrich the inner regions than the outer ones (e.g. Samland etĀ al., 1997; Prantzos & Boissier, 2000; DavĆ© etĀ al., 2011; Pilkington etĀ al., 2012; Gibson etĀ al., 2013; Hemler etĀ al., 2021; Tissera etĀ al., 2022). Flattened gradients could arise from radial mixing of gas and redistribution of metals on galaxy scales, induced by SN winds within the galaxy or galactic outflows of metal-enriched material expelled from the galaxy and re-accreted in the outer regions (so-called āgalactic fountainsā), driven by intense stellar feedback (e.g. Gibson etĀ al., 2013; Ma etĀ al., 2017), as well as by galaxy merging and interactions (e.g. Rupke etĀ al., 2010b, a; Rich etĀ al., 2012; Torres-Flores etĀ al., 2014). Finally, positive (radially increasing; also called āinvertedā) gradients may be produced by accretion of external pristine gas towards the central regions of the galaxy (e.g. Ceverino etĀ al., 2016); this could also contribute to flatten the metallicity gradient, if not invert it. Particularly strong metal-loaded galactic outflows could also contribute to produce a positive gradient, by moving the metal-rich gas from the central starburst regions to the outskirts (e.g. Tissera etĀ al., 2022). A schematic cartoon that summarises the above framework is shown in Fig. 1.
In the local Universe, most spiral galaxies show negative metallicity gradients, with more metal-enriched gas in the inner than in the outer regions of the galaxy (e.g. Magrini etĀ al., 2009, 2017; Stanghellini & Haywood, 2010; Luck etĀ al., 2011 for the Milky Way; Zaritsky etĀ al., 1994; Magrini etĀ al., 2010; Kewley etĀ al., 2010; Bresolin, 2011; SĆ”nchez etĀ al., 2014; Berg etĀ al., 2015; Ho etĀ al., 2015; Belfiore etĀ al., 2017 for other local galaxies). By using metallicity diagnostics that trace past metal enrichment, primarily planetary nebulae (PNe; up to 5ā10 Gyr ago, i.e. 2) some studies have suggested that the Milky Way and nearby galaxies had flatter gradients in the past which grew progressively steeper with cosmic time (Stanghellini etĀ al., 2014; Stanghellini & Haywood, 2018; Magrini etĀ al., 2016). A dependence of gradient slope with mass has been observed in the local Universe, with more massive galaxies exhibiting progressively steeper gradients, while low-mass galaxies having almost flat gradients (Belfiore etĀ al., 2017). This relation between mass and metallicity gradient may be the result of an evolutionary sequence in mass, with more massive, more evolved galaxies having steeper gradients and low-mass, less evolved ones, analogues of high- galaxies, having flatter gradients (e.g. Maiolino & Mannucci, 2019).
Cosmological simulations of galaxy evolution make very different predictions on metallicity gradients at high redshift (up to 3), which can strongly vary depending on prescriptions on star formation and stellar feedback models and on their relative contribution to gas mixing as compared to mergers (e.g. Gibson etĀ al., 2013; Ma etĀ al., 2017; Hemler etĀ al., 2021; Tissera etĀ al., 2022). For example, according to the FIRE simulations (Hopkins etĀ al., 2014), mergers and rapid variations of metallicity gradients in high- galaxies, induced by starburst episodes which drive strong outflows and disrupt the gas disc, would flatten any negative gradients previously developed, especially in low-mass galaxies where feedback mechanisms are more efficient (Ma etĀ al., 2017). Instead, the TNG50 simulations (Pillepich etĀ al., 2019; Nelson etĀ al., 2019) which feature less bursty feedback, predict negative gradients at all redshifts, steeper (more negative) with increasing (Hemler etĀ al., 2021). The metallicity gradients resulting from the MUGS (ānormalā feedback; Stinson etĀ al., 2010) and MaGICC (āenhancedā feedback; Brook etĀ al., 2012) simulation suites are compared in Gibson etĀ al. (2013). The āenhancedā feedback (including pre-SN early stellar feedback from massive stars) re-distributes energy and re-cycled ISM material over large scales, through re-accretion in the outer parts of the galaxy of gas expelled via outflows. In this case, relatively flat and temporally invariant abundance gradients are predicted, in contrast to the steeper negative gradients at increasing emerging from the normal feedback scenario. Finally, the EAGLE simulations (Schaye etĀ al., 2015) find that the median gradient is zero (flat) at all redshifts, but with the scatter around the median increasing with due to individual galaxies transitioning between steep (either negative or positive) and flatter gradients (Tissera etĀ al., 2022). The authors report that this behaviour results from the higher frequency of major mergers at higher leading to episodes of enhanced accretion of low-metallicity gas which trigger intense star formation and ejection of metal-enriched gas.
Observations at the cosmic noon epoch (i.e., 1ā4) have found heterogeneous results, with both negative, positive, and flat metallicity gradients (e.g. Cresci etĀ al., 2010; Yuan etĀ al., 2011; Queyrel etĀ al., 2012; Swinbank etĀ al., 2012; Stott etĀ al., 2014; Troncoso etĀ al., 2014; Jones etĀ al., 2013, 2015; Leethochawalit etĀ al., 2016; Wuyts etĀ al., 2016; Wang etĀ al., 2017, 2022; Carton etĀ al., 2018; Fƶrster Schreiber etĀ al., 2018; Curti etĀ al., 2020a). Nevertheless, the majority of high- measurements is consistent with little or no cosmic evolution of metallicity gradients, which are found to be flat or only moderately negative or positive ( dexĀ kpc; see e.g. the compilation in Curti etĀ al. 2020a).
At higher redshift ( 7), Vallini etĀ al. (2024) also found flat gradients by making use of a physically motivated Bayesian model to derive metallicities from rest-frame far-infrared lines ([Oāiii] 88m and [Cāii] 158m) from Atacama Large Millimeter/submillimeter Array (ALMA) observations. However, as the authors point out as a caveat, there may be systematics in comparing their results with gradients of O/H abundance obtained with the more standard methods based on optical lines, since the adopted model does not take into account the likely enhancement of O/C at low metallicities (see e.g. Maiolino & Mannucci, 2019), which would result in the model returning higher values of metallicity. Therefore, while their results consistent with flat gradients at 7 are very significant, they require independent confirmation.
The James Webb Space Telescope (JWST) has recently opened up the possibility of measuring metallicity gradients in the first 1ā2 Gyr of the Universe by using rest-frame optical emission lines (Wang etĀ al., 2022; RodrĆguez Del Pino etĀ al., 2023; Arribas etĀ al., 2023) which are extensively used and calibrated at lower redshifts (e.g. Curti etĀ al., 2017, 2020b).
In this work, we study the spatially resolved gas-phase metallicity and investigate the shape of metallicity gradients at high redshift ( 6), by tracing the warm (Ā Ā Ā K) ionised gas emission. Specifically, we observed three high- systems with the Near-InfraRed Spectrograph (NIRSpec) on board JWST in its Integral Field Unit (IFU) mode with the low-spectral resolution PRISM/CLEAR (āā100) disperser-filter combination. These are A2744-YD4, part of the proto-cluster A2744-z7p9OD (Ā Ā 7.88), BDF-3299 (Ā Ā 7.11), and COSMOS24108 (Ā Ā 6.36). Basic information on the targets is given in Table 1. We trace gas-phase metallicity by making use of strong-line calibrators relying on rest-frame optical and near-UV emission lines. We also present the basic properties of each source, specifically stellar mass, obtained from SED fitting, and SFR, from H or H. The star formation history of these targets is studied in detail in a separate paper (Kohandel et al., in prep.).
Throughout this work, the reported wavelengths are in vacuum and we adopt a flat CDM cosmology with 67.7Ā kmĀ sĀ Mpc, 0.31, and 0.69 (Planck Collaboration etĀ al., 2020).
2 Description of observations, data reduction, and analysis
The observations employed in this work were carried out between October and December 2022 as part of the JWST GO program ID 1893 (PI: S. Carniani; Carniani etĀ al. 2021). The three sources were observed for 15.102 ks (4 h) on-source each (22.059 ks with overheads; 6 h) with 1.678 ks of background exposure time each (2.830 ks with overheads). The observations were performed with the PRISM/CLEAR disperser-filter combination and an eight-point dither āmediumā cycling pattern (step 0.5). The JWST NIRSpec 3 IFU PRISM observations simultaneously span the spectral range 0.6ā5.2Ā m with a spectral resolution ranging between 30ā330.
Target name | RA [hh:mm:ss.ss] | Dec [dd:mm:ss.ss] | a | Scale [kpc arcsec] |
A2744-YD4 | +00:14:24.90 | ā30:22:56.10 | 7.879 | 3.512b |
BDF-3299 | +22:28:12.31 | ā35:10:00.39 | 7.114 | 5.289 |
COSMOS24108 | +10:00:47.33 | +02:28:43.14 | 6.361 | 5.647 |
a
From this work.
- b
2.1 Data reduction
We retrieved the raw data from the MAST archive and we ran the three stages of the pipeline using version 1.11 with CRDS (calibration reference data system) context ājwst_1094.pmapā. First, at stage 1 ācalwebb_detector1ā, the pipeline applied the detector-level corrections (e.g. check for saturation, dark exposure subtraction, flagging of bad pixels and cosmic-ray persistences) and performed ramp fitting for individual exposures. We then calibrated the count rate images by executing stage 2 ācalwebb_spec2ā of the pipeline, which corrects for flat field and performs the wavelength calibration. The background was subtracted from each exposure during this stage by using the observations of the dedicated background for each target. We processed the background targets up to stage 2 of the pipeline, then we applied the background step for the science target. In the background step, the pipeline subtract the background exposure from each target exposure in the detector space. Finally, each calibrated exposure was combined in stage 3 ācalwebb_spec3ā by using ādrizzleā weighting and a spaxel size of 0.05ā³Ā to obtain the final cubes. During stage 3, we applied the outlier rejection step built in the pipeline and then a sigma clipping to remove any residual outliers in the final cube.
2.2 Data analysis
Name | Emission line ratio |
R | [Oāii] 3727,30/H |
R | [Oāiii] 5008/H |
OO | [Oāiii] 5008/[Oāii] 3727,30 |
0.47 R + 0.88 R | |
NeO | [Neāiii] 3870/[Oāii] 3727,30 |
RS | [Oāiii] 5008/HĀ + [Sāii] 6718,33/H |
OS | [Oāiii] 5008/HĀ / [Sāii] 6718,33/H |
In this section we describe the analysis of the NIRSpec R100 IFU data. In brief, we obtained the emission line fluxes from emission-line modelling of the spectra, which were used to infer the gas-phase metallicity. This was done on both an integrated basis (to get the integrated metallicity of each target) and a spatially resolved one, in this case both spaxel-by-spaxel (to produce maps) and in concentric radial annuli (to obtain radial gradients). From the integrated spectra, we also obtained the stellar mass () of each source from SED fitting and their star formation rate (SFR) from the emission-line modelling. We provide more details in the following.
2.2.1 Emission line fitting
The main goal of this work is to measure radial gas-phase metallicity gradients. To do so, we infer the oxygen abundance relative to hydrogen (12+log(O/H)), a proxy of gas-phase metallicity, by making use of the optical and near-UV strong-line diagnostic ratios reported in Table 2. We adopt the new diagnostic ratio = 0.47 R + 0.88 R, first introduced in Laseter etĀ al. (2024), in place of the more traditional R = ([Oāii] 3727,30 + [Oāiii] 4960,5008)/H, since the former is more suited for high- galaxies, while the latter is a projection mostly driven by low-metallicity local analogues. We adopt the best-fit polynomial calibrations from Curti etĀ al. (2017, 2020b), slightly revisited in Curti etĀ al. (2023a, b) to better probe the low-O/H regime, to infer the gas metallicity from the combination of the above ratios. We tested other gas-phase metallicity calibrations (e.g. Nakajima etĀ al., 2022; Sanders etĀ al., 2024). These gave 12+log(O/H) values similar to the Curti et al. ones within 0.1 dex, consistent with the uncertainties, and virtually no difference in the metallicity radial gradients.
The FWHM of the point spread function (PSF) of JWST depends on wavelength. Specifically, the PSF FWHM of NIRSpec IFU ranges from around 0.09 at 1 m up to around 0.16 at 4.5 m, and the variation with wavelength is stronger in the direction perpendicular to the IFU slices than along them (DāEugenio etĀ al., 2023). Since the aim of this work is to obtain spatially resolved metallicities from emission line ratios sometimes far in wavelength (see Table 2), we applied a wavelength-dependent spatial smoothing to the data cube prior to the line fitting with the aim of achieving the same spatial resolution at all wavelengths. We adopted the PSF FWHM curves from DāEugenio etĀ al. (2023), empirically calibrated by matching NIRSpec IFU observations with NIRCam ones of the same target reported in their Eqs. 3 and 4, for the along- and across-slice cases, respectively. The smoothing was done by convolving the cube with a wavelength-dependent 2D Gaussian kernel in order to obtain a smoothed cube having the same spatial resolution at each wavelength, specifically the PSF FWHM reported in DāEugenio etĀ al. (2023) at the wavelength of the highest-wavelength line of interest for the along-slice case (which has the largest FWHM among the two cases). This corresponds to FWHM 0.16 (at [Oāiii]) for A2744-YD4, 0.14 (at [Oāiii]) for BDF-3299, and 0.18 (at H-[Sāii]) for COSMOS24108. Specifically, the 2D Gaussian kernel to be used for the convolution, at a given wavelength , was defined as follows: FWHM() = FWHM() ā FWHM(), where FWHM is the model FWHM from DāEugenio etĀ al. (2023) and ādirā indicates the direction, either along or across the slices.
In general, we employ the original unsmoothed data cube for the flux maps of continuum and emission lines and for integrated measurements from circular apertures, while we adopt the smoothed data cube for the maps and radial profiles involving line ratios and metallicity, which would otherwise be affected by the wavelength-dependence of the PSF. In case of A2744-YD4, the highest- target of our sample, we adopt the smoothed data cube for every map, to obtain visually clearer maps as compared to the more noisy ones from the unsmoothed cube.
The data analysis consisted in modelling the rest-frame optical and near-UV emission lines available in the observed spectral range. The emission lines are spectrally unresolved in the PRISM spectra (the spectral resolution is 1250 kmĀ s at 3 m and 450 kmĀ s at 5 m), therefore a single Gaussian function per line was used. We fitted the [Oāiii]-H, the [Oāii]-[Neāiii]-H, and the H-[Nāii]-[Sāii] line complexes separately. The [Nāii] 6550,85 was included to allow for a better modelling of the H line profile, which was otherwise showing a marked residual in its redward wing. This was done only when fitting integrated spectra from circular apertures or concentric annuli for radial profiles, not in the lower-S/N case of spaxel-by-spaxel fitting, for which no residual wing indicative of [Nāii] was detected above the noise. For each line complex, the velocity was tied to be the same for all the lines, while we allowed the line width to vary to match the PRISM spectral resolution, given that all lines are spectrally unresolved as mentioned. We fixed the flux ratios [Oāiii] 5008/4960 and [Nāii] 6585/6550 to their theoretical value of 3 (Storey & Zeippen, 2000). We included an underlying first-order polynomial to model the continuum. We accounted for the Balmer break of the continuum at around 3645 Ć Ā rest-frame, when needed, by employing three (first-order) polynomials, one modelling the jump and the other two on each side of it.
The line fluxes used to infer the metallicity or the SFR were first corrected for dust extinction, when possible (see below), using the H/H ratio (for A2744-YD4 and BDF-3299) or H/H (for COSMOS24108), only when both lines had S/N3. We adopted a Calzetti etĀ al. (2000) reddening curve, with an = 4.05, suitable for high- low-metallicity (12+log(O/H) 8.5) star-forming galaxies at the wavelengths of interest here of 3000Ā Ć Ā (see e.g. Shivaei etĀ al. 2020). We assumed the theoretical extinction-free H/H and H/H ratios of 0.466 and 2.87, respectively, valid for case-B recombination and an electron temperature of 10Ā K (Osterbrock & Ferland, 2006), typical of the warm ionised gas emitting rest-frame optical lines.
We selected the line ratio diagnostics for the metallicity estimation in an adaptive way, based on the line S/N. Specifically, a certain ratio (e.g. [Neāiii]/[Oāii]) was employed only when all the lines involved in it had S/N3. In this way, for each fitted spectrum we only selected the sub-sample of line ratio diagnostics whose line fluxes exceeded the S/N threshold, instead of using all the ratios (even those involving S/N3 lines) or none of them (when just one or a few of the lines were below the S/N threshold). Finally, given a set of diagnostic ratios, the best-fit metallicity was obtained by minimising the chi-squared defined simultaneously by the different observed ratios and their relative calibration curves, weighted by the observed uncertainties and the intrinsic dispersion of the calibration (added in quadrature; see Curti etĀ al. 2020b).
We first modelled the emission lines in integrated spectra extracted from circular apertures centred on each sub-source present in the three studied systems. From these, we obtained the integrated gas-phase metallicity (as described above) and SFR of each sub-source. The SFR was obtained from the extinction-corrected flux of H, using the relation from Kennicutt & Evans (2012). In A2744-YD4 and BDF-3299, for which H was not available, the SFR was obtained from the extinction-corrected flux of H, converted to H by adopting the theoretical H/H ratio of 2.87. We employed circular apertures of 0.15ā³Ā radius for A2744-YD4 and COSMOS24108, and of 0.1ā³Ā for BDF-3299, due to the spatial vicinity of the sources in this latter system.
The emission line modelling was then performed both on a spaxel-by-spaxel basis (for both the unsmoothed and the smoothed data cubes) and in concentric radial annuli centred on each source (for the smoothed data cube). The latter was done to increase the S/N on the emission lines and get more robust estimates of metallicity with the goal of obtaining metallicity gradients. For this annular line modelling, we extracted integrated spectra at each radius by collapsing the spaxels within concentric circular annuli. The annuli, having radial width of 1 spaxel (0.05) each, were centred on each of the main spatial components detected in each targeted system, with a variable maximum aperture radius depending on the component extension (as described in the next sections separately for each target).
The extinction-correction of emission line fluxes used to infer metallicity was done only for the case of integrated spectra from circular apertures and from concentric radial annuli. This was not possible for the spaxel-by-spaxel case because, among the useful Balmer lines tracing extinction, H is detected in almost no spaxels in A2744-YD4 and BDF-3299. Therefore, the maps of metallicity were obtained without accounting for possible extinction. For COSMOS24108, also including H in the spectral range, H is detected in a large enough number of spaxels to attempt for an extinction correction of the line fluxes for the metallicity. However, the resulting map is very similar to the non-extinction-corrected one, only more noisy, therefore we report the metallicity map obtained without the extinction correction.
Source | SFR | 12+log(O/H) | ||||||||
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) |
A2744-YD4 | ||||||||||
YD4 | 7.879 | 8.69 | 723 | 21 | 363 | 1804 | ā | ā | 2.2 | 8.19 |
YD6 | 7.880 | 8.40 | 263 | 20 | 17 | 443 | ā | ā | 1.0 | 8.27 |
YD6-[Oāiii]-E | 7.881 | 7.83 | 24 | 24 | 14 | 573 | ā | ā | 0.87 | |
YD4-[Oāiii]-W | 7.877 | 7.82 | 14 | 15 | 19 | 46 | ā | ā | 1.1 | |
ZD1 | 7.872 | 7.58 | 13 | 13 | 18 | 273 | ā | ā | 1.1 | |
YD1 | 7.883 | 8.58 | 1074 | 24 | 473 | 3463 | ā | ā | 2.85 | 8.11 |
YD1-E | 7.882 | 8.44 | 543 | 20 | 342 | 2702 | ā | ā | 2.02 | 7.98 |
s1 | 7.878 | 8.01 | 24 | 25 | 183 | 1513 | ā | ā | 1.09 | 7.72 |
BDF-3299 | ||||||||||
a | 7.114 | 7.90 | 242 | 13 | 361 | 1953 | ā | ā | 3.47 | 7.68 |
b | 7.114 | 8.21 | 282 | 11 | 442 | 2854 | ā | ā | 4.2 | 7.69 |
c | 7.112 | 7.62 | 13 | 12 | 152 | 934 | ā | ā | 1.5 | 7.80 |
COSMOS24108 | ||||||||||
a | 6.361 | 9.29 | 1795 | 486 | 535 | 3585 | 2194 | 344 | 14.4 | 8.20 |
b | 6.358 | 8.81 | 1506 | 476 | 724 | 4704 | 2453 | 18 | 9.8 | 8.11 |
c | 6.359 | 8.89 | 675 | 36 | 284 | 924 | 684 | 22 | 1.74 | 8.33 |
[Oāiii]-Ea | 6.362 | 8.53 | 623 | 363 | 413 | 2963 | 1833 | 21 | 14.5 | 8.01 |
[Oāiii]-Eb | 6.356 | 8.40 | 444 | 24 | 373 | 2443 | 1133 | 22 | 3.43 | 7.95 |
2.2.2 SED fitting
With the goal of obtaining the stellar mass of each target, we fitted the spatially integrated spectra extracted from the data cube by using the spectral energy distribution (SED) fitting code bagpipes (Carnall etĀ al., 2018). We adopted the stellar population models by Bruzual & Charlot (2003) and included the nebular emission with cloudy (Ferland etĀ al., 2017) with the ionisation parameter () as a free parameter and adopting a . We assumed a Kroupa (2001) initial mass function truncated at 0.01 and 100 M and a Calzetti etĀ al. (2000) attenuation curve. Finally, we used a non-parametric star-formation history model with continuity priors (see Leja etĀ al. 2019) and with four time bins: Myr, Myr, Myr, and Myr.
3 A2744-YD4
A2744-YD4 is part of the proto-cluster A2744-z7p9OD at Ā Ā 7.883 located behind the strong lensing cluster Abell 2744. The proto-cluster includes 22 sources at 7 Ā” Ā” 9 identified through combined deep HST and Spitzer IRAC photometry (Laporte etĀ al., 2014; Zheng etĀ al., 2014; Atek etĀ al., 2015; Ishigaki etĀ al., 2016) as part of the Hubble Frontier Fields program (Lotz etĀ al., 2017). Seven of them were recently spectroscopically confirmed to be at = 7.88 through JWST NIRSpec MSA spectroscopy (Morishita etĀ al., 2023). This makes A2744-z7p9OD the most distant proto-cluster known so far. The system is highly over-dense, with an excess of surface number density from the field average = 130 (Ishigaki etĀ al., 2016).
The 3 IFU FOV of our observations includes five sources of the proto-cluster known as the āquintetā, namely YD1, YD4, YD6, YD7, and ZD1, as well as a new source identified by Hashimoto etĀ al. (2023) and named s1 (Fig. 2, top-left). We only partially cover the East component of YD7 (named YD7-E), while its emission toward to West lies outside the FOV. The lensing magnification factor of the source in the FOV is 2.0 (Morishita etĀ al., 2023; Bergamini etĀ al., 2023). YD4, at the centre of our IFU observation, was reported to be at = 8.38 based on Ly and ALMA [Cāii]Ā 158 m and [Oāiii]Ā 88 m line emission (Laporte etĀ al., 2017, 2019; Carniani etĀ al., 2020). Recently, Morishita etĀ al. (2023) measured = 7.88 for YD4 based on the detection of high-S/N [Oāiii] 4959,5007 and H emission lines in high-spectral resolution (2700) NIRSpec MSA observations, thus ruling out the previous redshift measurement.
The top-left panel of Fig.Ā 2 shows the map of the continuum emission in the range 2ā3 m observed wavelength (0.23ā0.34Ā m rest-frame). In addition to the sources named YD4, YD6, YD1, s1, YD7-E, and ZD1, the continuum map also shows the presence of a source at the northernmost part of the FOV, which is also present in the NIRCam image but is not labelled in Hashimoto etĀ al. (2023); we name this source s2. However, given that this is at the edge of the FOV, where many artefacts are present in the NIRSpec IFU data, we consider this source as tentative and report its label in parenthesis in the figure. Moreover, in the map of the continuum obtained from the unsmoothed data cube integrated over the spectral range 1.2ā2 m (Fig. 10, top-left), where the spatial resolution is the highest, the eastern tail of YD1 appears as a separate peaked spatial component. We label this extra source as YD1-E.
Fig. 2, bottom panel, illustrates the integrated spectrum (aperture radius = 0.15) associated with the main component of the A2744-YD4 system, that is, YD4 itself. The rest-frame optical lines [Oāii], H, and [Oāiii] are detected with high S/N (6 on the peak). The spectroscopic redshift based on Lyman break and optical lines ( 7.88) is consistent with that found by Morishita etĀ al. (2023) and Hashimoto etĀ al. (2023) from the NIRSpec 2700 data. We find a stellar mass of log() 8.7 from the SED fitting and a SFR 2 Ā yr from H for YD4 (Table 3). These were corrected for the magnification factor of 2 due to the lensing. The gas-phase metallicity, relying on line ratios, is instead not affected by it. For the other spatial components in the system, we find log() 7.5ā8.6 and SFR 1ā3 Ā yr.
The [Oāiii] flux map resulting from our spectral emission-line modelling is displayed in Fig. 2, central panel (the maps of H and [Oāii] are shown in Fig. 10). From the [Oāiii] map, we identify two new emission regions which are weak or absent in continuum, which we label as YD6-[OIII]-E and YD4-[OIII]-W. Overall, the [Oāiii]Ā maps obtained from the PRISM data presented here are deeper and reveal fainter and more extended features than those from high-resolution grating data presented in Hashimoto etĀ al. (2023) (who do not report the detection of any other emission line due to the lower S/N of their data). We can count nine sources in total in the 3 NIRSpec IFU FOV, among those detected in continuum and in line emission, namely YD4, YD6, YD6-[OIII]-E, YD4-[OIII]-W, YD7-E, ZD1, YD1, YD1-E, and s1, and possibly a tenth source, s2. All these sources lie at 7.88, either confirmed spectroscopically (YD4, YD6, YD6-[OIII]-E, YD4-[OIII]-W, YD1, YD1-E, and s1; e.g. Morishita etĀ al. 2023, Hashimoto etĀ al. 2023, and this work) or from photometry (ZD1), except for YD7-E (and the tentative source s2) whose redshift is not assessed. The spectra of all targets, but YD4, are reported in Figs.Ā 8 and 9.
In the top-right panel of Fig. 2 we show the map of oxygen abundance, 12 + ā(O/H), a proxy for gas-phase metallicity. The metallicity is inferred by making use of the strong-line diagnostics reported in Table 2 as described in Sect. 2.2 (the maps of the [Oāiii] 5008/HĀ and [Oāiii] 5008/[Oāii] 3727,30Ā emission line ratios are shown in Fig. 10). The ratios involving lines which are not close in wavelength (all but [Oāiii]Ā and H) may be affected by extinction. Unfortunately, the only robustly detected Balmer line at the spaxel level is H, and only in a handful of spaxels, therefore estimating the dust extinction spaxel-by-spaxel is not possible. Based on this, the abundance map should not be taken as a robust measurement; moreover, the map is very noisy. Nevertheless, the map seems to suggest that, in the northern system, YD1 is embedded between two sources at lower metallicity, s1 and YD1-E. No clear pattern of metallicity shows up in the southern system (the YD4 one) from the map.
4 BDF-3299
BDF-3299 is a spectroscopically confirmed star-forming galaxy (SFR 6 Ā yr, SFR 12 /yr) at 7.109 (Castellano etĀ al., 2010; Vanzella etĀ al., 2011; Maiolino etĀ al., 2015; Carniani etĀ al., 2017), located in an over-density of galaxies (Castellano etĀ al., 2016). We report the maps for BDF-3299 from the NIRSpec IFU observations in Fig. 3.
The median continuum in the observed spectral range 1.1ā2 m (0.14ā0.25 m rest-frame) shows that BDF-3299 is composed of three spatial components. The two brighter ones are located at the NW (the strongest of the two in the continuum) and at the centre of the system, respectively; a third, fainter one resides in the E part. We label these sources as BDF-3299-a, BDF-3299-b, and BDF-3299-c, respectively, in decreasing order of continuum brightness, as labelled in Fig. 3, left panel. The other continuum emitters in the FOV are lower-redshift sources.
Fig. 3, bottom panel, shows the integrated spectrum (aperture radius = 0.1) associated with the brightest component in continuum emission, BDF-3299-a. The rest-frame optical lines [Oāii], H, and [Oāiii] are detected with high S/N (6 on the peak). From the lines, we estimate a redshift of 7.114, slightly higher but roughly consistent with the previously reported one from Ly (7.109; Vanzella etĀ al., 2011) and [Cāii] (7.107; Carniani etĀ al., 2017). We obtain a stellar mass of log() 7.9 from the SED fitting and a SFR 3.5 Ā yr from H for BDF-3299-a (Table 3). The spectra of the other spatial components in the system, namely BDF-3299-b and c, are reported in Fig.Ā 11. From these, we find log() 7.6ā8.2 and SFR 1.5ā4 Ā yr.
The [Oāiii] emission (Fig. 3, central panel) peaks on BDF-3299-b rather than on BDF-3299-a (the strongest in continuum emission). The line ratio maps are quite noisy and only few spaxels are above the S/N threshold of 3 (Fig. 12). Therefore, the same applies to the metallicity map, which generally shows low values, of 12+log(O/H) 7.8, in the few spaxels where it can be estimated (Fig. 3, right).
5 COSMOS-24108
COSMOS-24108 (Ā Ā 6.36; SFR 29 /yr, SFR 6.2 /yr) shows two, or possibly three, spatial components in H-band HST rest-frame UV imaging (Pentericci etĀ al., 2016). In Fig. 4, left panel, we show the 1ā2 m observed continuum (0.14ā0.27 m rest-frame) from the NIRSpec IFU data. This shows two main spatial components, and a third, fainter one in the SE part of the system, consistent with those seen with HST in a similar spectral band. We label these three spatial components as COSMOS24108-a, b, and c, in order of continuum brightness. The continuum at redder wavelengths, between 3ā3.5 m (0.39ā0.46 m rest-frame; Fig. 14), is dominated by the southernmost of the two main components, and only an extended emission towards the northernmost source is present at these wavelengths, instead of a more clearly separate spatial component as at lower wavelengths. Therefore, the northernmost source has bluer continuum than the southernmost one. In the SW corner of the FOV, another galaxy at a lower redshift is also present.
Fig. 4, bottom panel, shows the integrated spectrum (aperture radius = 0.15) associated with the brightest component in continuum emission, COSMOS24108-a. The rest-frame optical lines [Oāii], H, [Oāiii], H, and [Sāii] are detected with high S/N (4 on the peak). We obtain a stellar mass of log() 9.3 from SED fitting and a SFR 15 Ā yr from H for COSMOS24108-a (Table 3). The integrated spectra of the other spatial components in the system are shown in Fig.Ā 13. For these, we find log() 8.4ā8.9 and SFR 2ā15 Ā yr.
The [Oāiii] ionised gas line emission (Fig. 4, central panel; H, [Oāii], and H maps are reported in Fig. 14) is much more extended than the continuum, revealing two additional bright clumps to the SE of the system, one of the two also tentatively detected in continuum (see contours). We label these as COSMOS24108-[OIII]-Ea and [OIII]-Eb. The HĀ map (Fig. 14, bottom-left) shows a bridge of gas (visible also in [Oāiii], though weaker relative the rest of the emission) connecting the main system to the closer of these two clumps, COSMOS24108-[OIII]-Ea. By comparing the ionised gas line emission with the 1ā2 m continuum, we see that the former peaks in between the two main continuum components, with a preferential extension towards the northernmost, weaker continuum component (COSMOS24108-b) rather than to the southernmost, brighter one (COSMOS24108-a).
Fig.Ā 4, right panel, reports the map of gas metallicity (the [Oāiii] 5008/H and [Oāiii] 5008/[Oāii] 3727,30 line ratio maps are shown in Fig. 14). The metallicity exhibits differences of up to 0.5ā0.6 dex among the different sources in the system. The northernmost of the two main continuum sources, COSMOS24108-b, appears to have lower metallicity (12 + ā(O/H) 8.0ā8.1) than the southernmost one, COSMOS24108-a ( 8.2), while the SE minor continuum component, COSMOS24108-c, has larger metallicity ( 8.4). Towards the two ionised gas clumps to the SE, COSMOS24108-[OIII]-Ea and COSMOS24108-[OIII]-Eb, the metallicity is the lowest, with values of 7.8ā7.9.
6 Integrated mass-metallicity relation
We obtained the properties of each sub-source belonging to each target, by extracting integrated spectra from circular apertures of radius of 0.15 for A2744-YD4 and COSMOS24108 and of 0.1 for BDF-3299 centred on each sub-source (Table 3). We find emission line ratios of log([Oāiii]/H) 0.4ā0.9 for A2744-YD4, log([Oāiii]/H) 0.7ā0.8 for BDF-3299, and log([Oāiii]/H) 0.5ā0.9 and log([Sāii]/H) ā0.8 for COSMOS24108. These are consistent with the range of values found for 6ā8 star-forming galaxies in the JADES survey (Cameron etĀ al., 2023), of log([Oāiii]/H) 0.5ā0.8 and log([Sāii]/H) ā0.8 (when only considering detections and not upper limits), in a range of stellar masses (log() 6.5ā9.0) and SFRs (0.1ā30) comprising those of the sources in this work. In some cases, like COSMOS24108-a with log([Oāiii]/H) 0.8 and log([Sāii]/H) ā0.8, the ratios are at the high end of the values reported in (Cameron etĀ al., 2023).
We inferred the gas-phase metallicity and the SFR from emission-line fitting and the stellar mass from SED fitting, as described in Sect. 2.2. In Fig.Ā 5 we display the metallicity versus stellar mass diagram for the spatial components identified in the images (TableĀ 3). We compare these with the values reported by Curti etĀ al. (2023b) for the 3ā10 galaxies from the JADES (including GNz-11; Bunker etĀ al. 2023), CEERS (Nakajima etĀ al., 2023, re-computed by Curti etĀ al. 2023b for consistency), and EROs (Curti etĀ al., 2023a; Laseter etĀ al., 2024) samples (grey circles). The best-fit relation for this 3ā10 compilation obtained by Curti etĀ al. (2023b) is also shown. We also report for reference the best-fit mass-metallicity relations (MZRs) for galaxies at 0.08 from SDSS (Curti etĀ al., 2020b), 2ā3 from MOSDEF (Sanders etĀ al., 2021), and 3 from GLASS (Li etĀ al., 2023).
Our targets are generally compatible with the values found for the 3ā10 galaxies, especially the sources in BDF-3299 which sit on the MZR defined by this high- collection. We note that most of the sources in A2744-YD4 and COSMOS24108 ( 8 and 6, respectively) are at the high end of the JADES, CEERS, and EROs points, where the 2ā3 MZRs from MOSDEF and GLASS lie. This suggests that the A2744-YD4 and COSMOS24108 systems may comprise more evolved sources as compared to the majority of sources at the same redshifts and be instead more similar to galaxies at cosmic noon.
We further display the predictions for the high- MZR from a number of cosmological simulations, specifically EAGLE ( 8; Schaye etĀ al., 2015), FIRE ( 6; Ma etĀ al., 2016), IllustrisTNG ( 8; Torrey etĀ al., 2019), FirstLight ( 6 and 8; Langan etĀ al., 2020), and SERRA ( 8; Pallottini etĀ al., 2022)111The MZR from the SERRA simulation (Pallottini etĀ al., 2022) was originally presented in Curti etĀ al. (2023a) and was computed using 202 simulated galaxies at . In the present work, the MZR from SERRA is recomputed from an extended sample of 245 objects Pallottini & Ferrara (2023). This causes some differences at the low-mass end of the reported MZR with respect to that in Curti etĀ al. (2023a).. We note that the sources analysed in this work, both as a whole and within the individual A2744-YD4 and COSMOS24108 systems, are better aligned with simulations predicting steeper MZR slopes than the median of the observations at 3ā10.
7 Gas-phase metallicity gradients
Source name | [dex kpc] |
A2744-YD4 | 0.14 |
A2744-YD1 | ā0.05 |
BDF-3299 | ā0.03 |
COSMOS24108 | 0.03 |
To more robustly constrain the spatially resolved metallicity, we extracted the line fluxes from spatially integrated spectra from concentric radial annuli, as described in Sect. 2.2. This allowed us to increase the S/N of the emission lines at large radii and determine the dust extinction from the faint Balmer lines. We thus determined the gas-phase metallicity as a function of the distance from the centre of each source, defined as explained in the following.
The resulting metallicity gradients are reported in Fig. 6. In the system A2744-YD4, we determined the metallicity gradients only for the two brightest sources YD4 and YD1 (centred around their continuum peak) since the metallicity measurements for the other sources suffered from large uncertainties due to low S/N of emission lines. For BDF-3299, since we could have only extracted two radial points for each spatial component (BDF-3299-a, b, and c) due to the compactness of this system, we obtained the metallicity gradient profile for the whole BDF-3299 system, centred on the brightest source in line emission, BDF-3299-b. For the COSMOS24108 system, given that the line emission peaks between the two main continuum components, COSMOS24108-a and COSMOS24108-b, and not on either of them (see Fig. 4), we opted to extract a single set of concentric annular apertures centred on the line emission peak, rather than two sets centred on each of the continuum sources (since we considered this latter approach not very sensible in this case). We also report the metallicity gradient from two additional smaller apertures centred on the emission line clumps to the SE of the main system, COSMOS24108-Ea and COSMOS24108-Eb.
For all the sources, A2744-YD4, YD1, BDF-3299, and COSMOS24108, the metallicity gradients span a range between and (TableĀ 4). We note that, within the uncertainties, all metallicity gradients in our sample are flat.
8 Discussion
We found flat gradients within the uncertainties for the sources in A2744-YD4, BDF-3299, and COSMOS24108 analysed in this work. As introduced in Sect. 1 and sketched in Fig.Ā 1, negative (radially decreasing) gradients are expected as a consequence of the inside-out galaxy formation scenario, while inverted (radially increasing) gradients may arise as a consequence of pristine gas accretion towards the central regions. Possible explanations for flat radial gradients are radial gas mixing processes, occurring as a consequence of mergers, SN-driven winds and/or large-scale gas circulation (āgalactic fountainsā, i.e. metal-loaded outflows expelled from the galaxy and re-accreted in the external regions), which redistribute the gas in the galaxy and wash out any pre-existing radial gradient of metallicity.
All the targets in this work show multiple sources within a few kpc and disturbed morphologies. This indicates that they are experiencing, or have experienced, interactions and galaxy merging processes, which are expected to be more frequent in the early Universe according to cosmological simulations (see e.g. Kohandel etĀ al., 2020; Pallottini & Ferrara, 2023). Specifically, galaxies at 6 with stellar masses 510 living in dense environments, similar to those studied in this work (see TableĀ 3), are expected to have already experienced multiple merger events (Gelli etĀ al., 2020). On the other hand, while in the low-spectral resolution ( 100) NIRSpec PRISM data analysed in this work the emission lines are spectrally unresolved, [Oāiii] and H do not show evidence for asymmetric wings or any complex line profile indicative of outflows in the higher resolution ( 2700) NIRSpec IFU or MSA data from Hashimoto etĀ al. (2023) and Morishita etĀ al. (2023), respectively, for the case of YD4, YD1, and s1 in the A2744-YD4 system (YD4 having both IFU and MSA 2700 data). Therefore, this indicates that strong outflows are not occurring at present in this system. However, in principle we cannot exclude that the redistribution of metals across the galaxy may also be the result of past intense SN winds mixing the ISM or re-accretion in the outer regions of metal-enriched gas expelled by past galactic outflows, whose trace is absent in present-day spectra. All in all, mergers and/or possibly either past galactic outflows or SN wind mixing seem the most likely mechanism driving the observed flat metallicity gradients.
We stress that, even if we did our best in extracting the metallicity gradient for each separate spatial component, complications arise because in some cases these sub-sources are very close to each other and may contaminate each otherās metallicity (as appears to be the case of YD1 and YD1-E; Fig. 2) or there is confusion as to which source the detected gas belongs (as in the case of COSMOS24108-a and COSMOS24108-b, where gas is located between the two sources). Moreover, the sources are not settled in radially symmetric metallicity distributions. All these aspects can contribute to yielding a flat gradient. Nevertheless, the spatial vicinity of the sources and the displacement of gas and stars are the result of the ongoing merging processes. Therefore, ultimately, mergers appear to be the most likely cause for the flatness of the metallicity gradients.
8.1 Metallicity gradients across cosmic time
In Fig. 7, we report the metallicity gradients from a compilation of observational studies, from = 0 up to 8 (this study), namely the highest redshift probed so far. The flat gradients within the uncertainties that we find at 6ā8 are compatible with the other currently available measured gradients at these redshifts, namely Vallini etĀ al. (2024) from FIR lines with ALMA and Arribas etĀ al. (2023) from rest-frame optical lines with JWST/NIRSpec. Specifically, the results from Vallini etĀ al. (2024) showcase the potential for synergies between spatially resolved observations with JWST and ALMA, though the spatial resolutions may be quite different ( 0.2ā0.4 kpc in our case versus 1.5 kpc in the case of Vallini etĀ al. 2024).
We also report the predictions for the evolution of metallicity gradients with redshift from cosmological simulations. Some of them predict steeper negative gradients with increasing redshift (MUGS, with normal feedback, Gibson etĀ al. 2013; TNG50, Hemler etĀ al. 2021), while others predict median flatter gradients at higher though with increasing scatter (MAGICC, with enhanced feedback, Gibson etĀ al. 2013; FIRE, Ma etĀ al. 2017; EAGLE, Tissera etĀ al. 2022).
Unfortunately, predictions for the cosmic evolution of metallicity gradients from simulations are limited to 3 so far. Nevertheless, if we assume that the extrapolation at higher follows the same trend which is in place up to 3, we can compare the predictions with our measurements at 6ā8. Specifically, a rough linear extrapolation of the trends from TNG50 (Hemler etĀ al., 2021) and MUGS (Gibson etĀ al., 2013) would give a metallicity gradient at 8 of ā0.2 dexĀ kpc and ā0.5 dexĀ kpc, respectively, which are both incompatible with our estimates considering the uncertainties (Fig. 7).
Our results, showing flat gradients within the uncertainties, are more compatible with expectations from those cosmological simulations predicting flatter gradients at high (MaGICC, Gibson etĀ al. 2013; FIRE, Ma etĀ al. 2017; EAGLE, Tissera etĀ al. 2022). This conclusion is based on the extrapolation of current model predictions, therefore we stress the need for updated metallicity gradient predictions from simulations at 3 in order to match the redshifts reached by current observations.
We point out that the merger fraction increases with redshift (e.g. theoretically Fakhouri etĀ al. 2010; Rodriguez-Gomez etĀ al. 2015; OāLeary etĀ al. 2021; observationally Duncan etĀ al. 2019), therefore galaxies at high redshift often have disturbed morphologies due to interactions and do not have radially symmetric gas and chemical distributions (see e.g. Figs. 2, 3, and 4; also e.g. observations in de Graaff etĀ al. 2023; Arribas etĀ al. 2023; RodrĆguez Del Pino etĀ al. 2023 and simulations in Pallottini etĀ al. 2017, 2019). Specifically, the metallicity maps in our work show in some cases non-azimuthally symmetric structures which are averaged out when producing radial gradients (as it is e.g. the case of YD1, having the highest values to the W and the lowest to the E of its spatial emission peak). Therefore, radial gradients may not be the optimal means to describe the distribution of metals in high- galaxies and other alternative methods to quantify it (such as the metallicity scatter in the galaxy, as suggested by Maiolino & Mannucci 2019) should be sought and included in predictions from simulations.
In summary, the flat radial gradients and the asymmetric spatial distribution of metallicity observed in some cases seem to support the merging scenario, though ISM mixing due to past SN winds or re-accretion in the outer regions of metal-enriched gas from past galactic outflows (galactic fountains) are also a viable mechanism to explain the flatness of the gradients. Spatially asymmetric low-metallicity accretion from the CGM and IGM could also contribute to produce non-azimuthally symmetric metallicities as suggested by Arribas etĀ al. (2023).
9 Conclusions
In this work, we have presented new JWST NIRSpec IFU observations at low spectral resolution (āā100; PRISM/CLEAR) of three high- systems, namely A2744-YD4 in the proto-cluster A2744-z7p9OD (Ā Ā 7.88), BDF-3299 (Ā Ā 7.11), and COSMOS24108 (Ā Ā 6.36). At these redshifts, the NIRSpec PRISM spectra (spanning 0.6ā5.2 m) cover the rest-frame UV and optical spectral ranges and include the main optical emission lines from warm (Ā Ā Ā K) ionised gas, [Oāii], H, H, [Oāiii], and, for COSMOS24108, also H and [Sāii]. We modelled the ionised gas lines in the spectra with Gaussian functions and mapped their emission. The main goal of this work is to study the spatially resolved gas-phase metallicity and investigate the shape of metallicity gradients at high redshift ( 6).
The targets have very disturbed morphologies in both the stellar continuum and ionised gas line emission. We identify several spatial components concentrated within a few kiloparsecs in all the three systems. We obtained the main integrated properties of each spatial component in each target, specifically stellar mass from SED fitting, SFR from H or H, and gas-phase metallicity by means of flux ratios of rest-frame optical emission lines which we modelled with Gaussian functions. We found log() 7.6ā9.3, SFRs 1ā15 Ā yr, and gas-phase metallicities 12+log(O/H) 7.7ā8.3 by extracting integrated spectra from circular apertures centred on each spatial component.
We compared the stellar masses and gas-phase metallicities of the targets with the mass-metallicity relations (MZRs) inferred observationally at different redshifts and with those predicted by cosmological simulations. In general, the sources in the systems studied in this work are consistent with the distribution of the highest- galaxies to date ( 3ā10) from JADES, CEERS, and EROs samples in the mass-metallicity plane. Nevertheless, most of the sources in A2744-YD4 and COSMOS24108 lie at the upper end of this 3ā10 distribution, in terms of metallicity at a given stellar mass, and are closer to the best-fit MZRs measured at 2ā3 (e.g. MOSDEF and GLASS surveys) than to the MZR at 3ā10. Relative to the MZRs from cosmological simulations, the sources in the three systems studied in this work are in good agreement with the slopes predicted at 6ā8 by most simulations.
We inferred the gas-phase metallicity radial gradients by extracting integrated spectra from concentric radial annuli centred on the main sources of each system for which the S/N allowed for a robust metallicity estimate. The gas-phase metallicity gradients are flat within the uncertainties. Flat gradients can be associated with processes which mix the gas and the metals on galaxy scales and therefore wash out any gradient which may have formed in the galaxy, such as mergers, SN wind mixing, and re-accretion of metal-loaded galactic outflows in the outer regions.
These are among the very few measurements of spatially resolved gas metallicity at these high redshifts ( 6ā8). In particular, YD4 (at 8) constitutes the highest- source in which a metallicity gradient has been probed so far.
Cosmological simulations make very different predictions regarding the cosmic evolution of metallicity gradients, though they are limited to 2ā3. Some of them predict steeper negative gradients with increasing redshift (e.g. MUGS and TNG50), while others predict median flatter gradients at higher (e.g. MaGICC, FIRE, and EAGLE). The results of this work, indicating mainly flat gradients, are more compatible with the predictions from this latter group of simulations, when considering their extrapolation at 6ā8.
All in all, the results of this work in terms of the MZR and gas-phase metallicity gradients at 6ā8 provide important constraints to guide future cosmological simulations. In particular, they urge for specific predictions on the cosmic evolution of metallicity gradients and metallicity maps out to redshift of 8, given that such predictions are currently limited to 3.
Moreover, galaxies at high tend to have more irregular morphologies and may not have azimuthally symmetric chemical distributions, as a result of frequent mergers and asymmetric low-mass gas infall. Radial gradients, which average out any azimuthal information, may then not be the optimal means to quantify the distribution of metals in high- galaxies. Therefore, other alternative quantitative tracers of the spatially resolved metallicity should be considered as part of predictions from simulations, to be compared with observations.
Acknowledgements.
We acknowledge support from European Unionās HE ERC Starting Grant No. 101040227 - WINGS (G.V., S.C.), INAF Minigrant āReionization and fundamental cosmology with high-redshift galaxiesā (M.C.), āFirstGalaxiesā Advanced Grant from the European Research Council (ERC) under the European Unionās Horizon 2020 research and innovation program, Grant agreement No. 789056 (A.J.B.), ERC Advanced Grant INTERSTELLAR H2020/740120 (A.F.), grant PID2021-127718NB-I00 funded by the Spanish Ministry of Science and Innovation/State Agency of Research, MICIN/AEI/ 10.13039/501100011033 (S.A.). This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1893. This research has made use of NASAās Astrophysics Data System Bibliographic Services. This research has made use of the Python packages NumPy (Harris etĀ al., 2020), SciPy (Virtanen etĀ al., 2020), IPython (Perez & Granger, 2007), Matplotlib (Hunter, 2007), emcee (Foreman-Mackey etĀ al., 2013), and Astropy (http://www.astropy.org), a community-developed core Python package for Astronomy (Astropy Collaboration etĀ al., 2013, 2018). This research has made use of the FITS files visualisation tool QFitsView (https://www.mpe.mpg.de/~ott/QFitsView/).References
- Arribas etĀ al. (2023) Arribas, S., Perna, M., RodrĆguez Del Pino, B., etĀ al. 2023, arXiv e-prints, arXiv:2312.00899
- Astropy Collaboration etĀ al. (2018) Astropy Collaboration, Price-Whelan, A.Ā M., SipÅcz, B.Ā M., etĀ al. 2018, AJ, 156, 123
- Astropy Collaboration etĀ al. (2013) Astropy Collaboration, Robitaille, T.Ā P., Tollerud, E.Ā J., etĀ al. 2013, A&A, 558, A33
- Atek etĀ al. (2015) Atek, H., Richard, J., Kneib, J.-P., etĀ al. 2015, ApJ, 800, 18
- Belfiore etĀ al. (2017) Belfiore, F., Maiolino, R., Tremonti, C., etĀ al. 2017, MNRAS, 469, 151
- Berg etĀ al. (2015) Berg, D.Ā A., Skillman, E.Ā D., Croxall, K.Ā V., etĀ al. 2015, ApJ, 806, 16
- Bergamini etĀ al. (2023) Bergamini, P., Acebron, A., Grillo, C., etĀ al. 2023, ApJ, 952, 84
- Bresolin (2011) Bresolin, F. 2011, ApJ, 730, 129
- Brook etĀ al. (2012) Brook, C.Ā B., Stinson, G., Gibson, B.Ā K., Wadsley, J., & Quinn, T. 2012, MNRAS, 424, 1275
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Bunker etĀ al. (2023) Bunker, A.Ā J., Saxena, A., Cameron, A.Ā J., etĀ al. 2023, A&A, 677, A88
- Calzetti etĀ al. (2000) Calzetti, D., Armus, L., Bohlin, R.Ā C., etĀ al. 2000, ApJ, 533, 682
- Cameron etĀ al. (2023) Cameron, A.Ā J., Saxena, A., Bunker, A.Ā J., etĀ al. 2023, A&A, 677, A115
- Carnall et al. (2018) Carnall, A. C., McLure, R. J., Dunlop, J. S., & Davé, R. 2018, MNRAS, 480, 4379
- Carniani etĀ al. (2021) Carniani, S., Arribas, S., Bunker, A., etĀ al. 2021, Galaxy assembly at z Āæ 6: unraveling the origin of the spatial offset between the UV and FIR emission, JWST Proposal. Cycle 1, ID. #1893
- Carniani etĀ al. (2020) Carniani, S., Ferrara, A., Maiolino, R., etĀ al. 2020, MNRAS, 499, 5136
- Carniani etĀ al. (2017) Carniani, S., Maiolino, R., Pallottini, A., etĀ al. 2017, A&A, 605, A42
- Carton etĀ al. (2018) Carton, D., Brinchmann, J., Contini, T., etĀ al. 2018, MNRAS, 478, 4293
- Castellano etĀ al. (2016) Castellano, M., Dayal, P., Pentericci, L., etĀ al. 2016, ApJ, 818, L3
- Castellano etĀ al. (2010) Castellano, M., Fontana, A., Paris, D., etĀ al. 2010, A&A, 524, A28
- Ceverino et al. (2016) Ceverino, D., SÔnchez Almeida, J., Muñoz Tuñón, C., et al. 2016, MNRAS, 457, 2605
- Cresci etĀ al. (2010) Cresci, G., Mannucci, F., Maiolino, R., etĀ al. 2010, Nature, 467, 811
- Curti etĀ al. (2017) Curti, M., Cresci, G., Mannucci, F., etĀ al. 2017, MNRAS, 465, 1384
- Curti etĀ al. (2023a) Curti, M., DāEugenio, F., Carniani, S., etĀ al. 2023a, MNRAS, 518, 425
- Curti etĀ al. (2020a) Curti, M., Maiolino, R., Cirasuolo, M., etĀ al. 2020a, MNRAS, 492, 821
- Curti etĀ al. (2023b) Curti, M., Maiolino, R., Curtis-Lake, E., etĀ al. 2023b, arXiv e-prints, arXiv:2304.08516
- Curti etĀ al. (2020b) Curti, M., Mannucci, F., Cresci, G., & Maiolino, R. 2020b, MNRAS, 491, 944
- Davé et al. (2011) Davé, R., Finlator, K., & Oppenheimer, B. D. 2011, MNRAS, 416, 1354
- de Graaff etĀ al. (2023) de Graaff, A., Rix, H.-W., Carniani, S., etĀ al. 2023, arXiv e-prints, arXiv:2308.09742
- DāEugenio etĀ al. (2023) DāEugenio, F., Perez-Gonzalez, P., Maiolino, R., etĀ al. 2023, arXiv e-prints, arXiv:2308.06317
- Duncan etĀ al. (2019) Duncan, K., Conselice, C.Ā J., Mundy, C., etĀ al. 2019, ApJ, 876, 110
- Fakhouri etĀ al. (2010) Fakhouri, O., Ma, C.-P., & Boylan-Kolchin, M. 2010, MNRAS, 406, 2267
- Ferland et al. (2017) Ferland, G. J., Chatzikos, M., GuzmÔn, F., et al. 2017, Rev. Mexicana Astron. Astrofis., 53, 385
- Foreman-Mackey etĀ al. (2013) Foreman-Mackey, D., Hogg, D.Ā W., Lang, D., & Goodman, J. 2013, PASP, 125, 306
- Förster Schreiber et al. (2018) Förster Schreiber, N. M., Renzini, A., Mancini, C., et al. 2018, ApJS, 238, 21
- Gelli etĀ al. (2020) Gelli, V., Salvadori, S., Pallottini, A., & Ferrara, A. 2020, MNRAS, 498, 4134
- Gibson etĀ al. (2013) Gibson, B.Ā K., Pilkington, K., Brook, C.Ā B., Stinson, G.Ā S., & Bailin, J. 2013, A&A, 554, A47
- Gillman etĀ al. (2021) Gillman, S., Tiley, A.Ā L., Swinbank, A.Ā M., etĀ al. 2021, MNRAS, 500, 4229
- Harris etĀ al. (2020) Harris, C.Ā R., Millman, K.Ā J., vanĀ der Walt, S.Ā J., etĀ al. 2020, Nature, 585, 357
- Hashimoto etĀ al. (2023) Hashimoto, T., Ćlvarez-MĆ”rquez, J., Fudamoto, Y., etĀ al. 2023, ApJ, 955, L2
- Hemler etĀ al. (2021) Hemler, Z.Ā S., Torrey, P., Qi, J., etĀ al. 2021, MNRAS, 506, 3024
- Ho etĀ al. (2015) Ho, I.Ā T., Kudritzki, R.-P., Kewley, L.Ā J., etĀ al. 2015, MNRAS, 448, 2030
- Hopkins et al. (2014) Hopkins, P. F., KereŔ, D., Oñorbe, J., et al. 2014, MNRAS, 445, 581
- Hunter (2007) Hunter, J.Ā D. 2007, Computing in Science and Engineering, 9, 90
- Ishigaki etĀ al. (2016) Ishigaki, M., Ouchi, M., & Harikane, Y. 2016, ApJ, 822, 5
- Jones etĀ al. (2013) Jones, T., Ellis, R.Ā S., Richard, J., & Jullo, E. 2013, ApJ, 765, 48
- Jones etĀ al. (2015) Jones, T., Wang, X., Schmidt, K.Ā B., etĀ al. 2015, AJ, 149, 107
- Kennicutt & Evans (2012) Kennicutt, R.Ā C. & Evans, N.Ā J. 2012, ARA&A, 50, 531
- Kewley etĀ al. (2010) Kewley, L.Ā J., Rupke, D., Zahid, H.Ā J., Geller, M.Ā J., & Barton, E.Ā J. 2010, ApJ, 721, L48
- Kohandel etĀ al. (2020) Kohandel, M., Pallottini, A., Ferrara, A., etĀ al. 2020, MNRAS, 499, 1250
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Langan etĀ al. (2020) Langan, I., Ceverino, D., & Finlator, K. 2020, MNRAS, 494, 1988
- Laporte etĀ al. (2017) Laporte, N., Ellis, R.Ā S., Boone, F., etĀ al. 2017, ApJ, 837, L21
- Laporte etĀ al. (2019) Laporte, N., Katz, H., Ellis, R.Ā S., etĀ al. 2019, MNRAS, 487, L81
- Laporte etĀ al. (2014) Laporte, N., Streblyanska, A., Clement, B., etĀ al. 2014, A&A, 562, L8
- Laseter etĀ al. (2024) Laseter, I.Ā H., Maseda, M.Ā V., Curti, M., etĀ al. 2024, A&A, 681, A70
- Leethochawalit etĀ al. (2016) Leethochawalit, N., Jones, T.Ā A., Ellis, R.Ā S., etĀ al. 2016, ApJ, 820, 84
- Leja etĀ al. (2019) Leja, J., Carnall, A.Ā C., Johnson, B.Ā D., Conroy, C., & Speagle, J.Ā S. 2019, ApJ, 876, 3
- Li etĀ al. (2023) Li, M., Cai, Z., Bian, F., etĀ al. 2023, ApJ, 955, L18
- Lotz etĀ al. (2017) Lotz, J.Ā M., Koekemoer, A., Coe, D., etĀ al. 2017, ApJ, 837, 97
- Luck etĀ al. (2011) Luck, R.Ā E., Andrievsky, S.Ā M., Kovtyukh, V.Ā V., Gieren, W., & Graczyk, D. 2011, AJ, 142, 51
- Ma etĀ al. (2016) Ma, X., Hopkins, P.Ā F., Faucher-GiguĆØre, C.-A., etĀ al. 2016, MNRAS, 456, 2140
- Ma etĀ al. (2017) Ma, X., Hopkins, P.Ā F., Feldmann, R., etĀ al. 2017, MNRAS, 466, 4780
- Magrini etĀ al. (2016) Magrini, L., Coccato, L., Stanghellini, L., Casasola, V., & Galli, D. 2016, A&A, 588, A91
- Magrini etĀ al. (2017) Magrini, L., Randich, S., Kordopatis, G., etĀ al. 2017, A&A, 603, A2
- Magrini etĀ al. (2009) Magrini, L., Sestito, P., Randich, S., & Galli, D. 2009, A&A, 494, 95
- Magrini etĀ al. (2010) Magrini, L., Stanghellini, L., Corbelli, E., Galli, D., & Villaver, E. 2010, A&A, 512, A63
- Maiolino etĀ al. (2015) Maiolino, R., Carniani, S., Fontana, A., etĀ al. 2015, MNRAS, 452, 54
- Maiolino & Mannucci (2019) Maiolino, R. & Mannucci, F. 2019, A&AĀ Rev., 27, 3
- Morishita etĀ al. (2023) Morishita, T., Roberts-Borsani, G., Treu, T., etĀ al. 2023, ApJ, 947, L24
- Nakajima etĀ al. (2023) Nakajima, K., Ouchi, M., Isobe, Y., etĀ al. 2023, ApJS, 269, 33
- Nakajima etĀ al. (2022) Nakajima, K., Ouchi, M., Xu, Y., etĀ al. 2022, ApJS, 262, 3
- Nelson etĀ al. (2019) Nelson, D., Pillepich, A., Springel, V., etĀ al. 2019, MNRAS, 490, 3234
- OāLeary etĀ al. (2021) OāLeary, J.Ā A., Moster, B.Ā P., Naab, T., & Somerville, R.Ā S. 2021, MNRAS, 501, 3215
- Osterbrock & Ferland (2006) Osterbrock, D.Ā E. & Ferland, G.Ā J. 2006, Astrophysics of gaseous nebulae and active galactic nuclei
- Pallottini & Ferrara (2023) Pallottini, A. & Ferrara, A. 2023, A&A, 677, L4
- Pallottini etĀ al. (2019) Pallottini, A., Ferrara, A., Decataldo, D., etĀ al. 2019, MNRAS, 487, 1689
- Pallottini etĀ al. (2022) Pallottini, A., Ferrara, A., Gallerani, S., etĀ al. 2022, MNRAS, 513, 5621
- Pallottini etĀ al. (2014) Pallottini, A., Ferrara, A., Gallerani, S., Salvadori, S., & DāOdorico, V. 2014, MNRAS, 440, 2498
- Pallottini etĀ al. (2017) Pallottini, A., Ferrara, A., Gallerani, S., etĀ al. 2017, MNRAS, 465, 2540
- Pentericci etĀ al. (2016) Pentericci, L., Carniani, S., Castellano, M., etĀ al. 2016, ApJ, 829, L11
- Perez & Granger (2007) Perez, F. & Granger, B.Ā E. 2007, Computing in Science and Engineering, 9, 21
- Pilkington etĀ al. (2012) Pilkington, K., Few, C.Ā G., Gibson, B.Ā K., etĀ al. 2012, A&A, 540, A56
- Pillepich etĀ al. (2019) Pillepich, A., Nelson, D., Springel, V., etĀ al. 2019, MNRAS, 490, 3196
- Planck Collaboration etĀ al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., etĀ al. 2020, A&A, 641, A6
- Popesso etĀ al. (2023) Popesso, P., Concas, A., Cresci, G., etĀ al. 2023, MNRAS, 519, 1526
- Prantzos & Boissier (2000) Prantzos, N. & Boissier, S. 2000, MNRAS, 313, 338
- Queyrel etĀ al. (2012) Queyrel, J., Contini, T., Kissler-Patig, M., etĀ al. 2012, A&A, 539, A93
- Rich etĀ al. (2012) Rich, J.Ā A., Torrey, P., Kewley, L.Ā J., Dopita, M.Ā A., & Rupke, D.Ā S.Ā N. 2012, ApJ, 753, 5
- RodrĆguez Del Pino etĀ al. (2023) RodrĆguez Del Pino, B., Perna, M., Arribas, S., etĀ al. 2023, arXiv e-prints, arXiv:2309.14431
- Rodriguez-Gomez etĀ al. (2015) Rodriguez-Gomez, V., Genel, S., Vogelsberger, M., etĀ al. 2015, MNRAS, 449, 49
- Rupke etĀ al. (2010a) Rupke, D., Kewley, L., & Barnes, J. 2010a, in Astronomical Society of the Pacific Conference Series, Vol. 423, Galaxy Wars: Stellar Populations and Star Formation in Interacting Galaxies, ed. B.Ā Smith, J.Ā Higdon, S.Ā Higdon, & N.Ā Bastian, 355
- Rupke etĀ al. (2010b) Rupke, D. S.Ā N., Kewley, L.Ā J., & Barnes, J.Ā E. 2010b, ApJ, 710, L156
- Rupke etĀ al. (2010c) Rupke, D. S.Ā N., Kewley, L.Ā J., & Chien, L.Ā H. 2010c, ApJ, 723, 1255
- Samland etĀ al. (1997) Samland, M., Hensler, G., & Theis, C. 1997, ApJ, 476, 544
- SÔnchez et al. (2014) SÔnchez, S. F., Rosales-Ortega, F. F., Iglesias-PÔramo, J., et al. 2014, A&A, 563, A49
- Sanders etĀ al. (2021) Sanders, R.Ā L., Shapley, A.Ā E., Jones, T., etĀ al. 2021, ApJ, 914, 19
- Sanders etĀ al. (2024) Sanders, R.Ā L., Shapley, A.Ā E., Topping, M.Ā W., Reddy, N.Ā A., & Brammer, G.Ā B. 2024, ApJ, 962, 24
- Schaye etĀ al. (2015) Schaye, J., Crain, R.Ā A., Bower, R.Ā G., etĀ al. 2015, MNRAS, 446, 521
- Shivaei etĀ al. (2020) Shivaei, I., Reddy, N., Rieke, G., etĀ al. 2020, ApJ, 899, 117
- Simons etĀ al. (2021) Simons, R.Ā C., Papovich, C., Momcheva, I., etĀ al. 2021, ApJ, 923, 203
- Somerville & Davé (2015) Somerville, R. S. & Davé, R. 2015, ARA&A, 53, 51
- Stanghellini & Haywood (2010) Stanghellini, L. & Haywood, M. 2010, ApJ, 714, 1096
- Stanghellini & Haywood (2018) Stanghellini, L. & Haywood, M. 2018, ApJ, 862, 45
- Stanghellini etĀ al. (2014) Stanghellini, L., Magrini, L., Casasola, V., & Villaver, E. 2014, A&A, 567, A88
- Stinson etĀ al. (2010) Stinson, G.Ā S., Bailin, J., Couchman, H., etĀ al. 2010, MNRAS, 408, 812
- Storey & Zeippen (2000) Storey, P.Ā J. & Zeippen, C.Ā J. 2000, MNRAS, 312, 813
- Stott etĀ al. (2014) Stott, J.Ā P., Sobral, D., Swinbank, A.Ā M., etĀ al. 2014, MNRAS, 443, 2695
- Swinbank etĀ al. (2012) Swinbank, A.Ā M., Sobral, D., Smail, I., etĀ al. 2012, MNRAS, 426, 935
- Tissera etĀ al. (2022) Tissera, P.Ā B., Rosas-Guevara, Y., Sillero, E., etĀ al. 2022, MNRAS, 511, 1667
- Torres-Flores etĀ al. (2014) Torres-Flores, S., Scarano, S., Mendes de Oliveira, C., etĀ al. 2014, MNRAS, 438, 1894
- Torrey etĀ al. (2019) Torrey, P., Vogelsberger, M., Marinacci, F., etĀ al. 2019, MNRAS, 484, 5587
- Troncoso etĀ al. (2014) Troncoso, P., Maiolino, R., Sommariva, V., etĀ al. 2014, A&A, 563, A58
- Vallini etĀ al. (2024) Vallini, L., Witstok, J., Sommovigo, L., etĀ al. 2024, MNRAS, 527, 10
- Vanzella etĀ al. (2011) Vanzella, E., Pentericci, L., Fontana, A., etĀ al. 2011, ApJ, 730, L35
- Virtanen etĀ al. (2020) Virtanen, P., Gommers, R., Oliphant, T.Ā E., etĀ al. 2020, Nature Methods, 17, 261
- Wang etĀ al. (2022) Wang, X., Jones, T., Vulcani, B., etĀ al. 2022, ApJ, 938, L16
- Wang etĀ al. (2020) Wang, X., Jones, T.Ā A., Treu, T., etĀ al. 2020, ApJ, 900, 183
- Wang etĀ al. (2019) Wang, X., Jones, T.Ā A., Treu, T., etĀ al. 2019, ApJ, 882, 94
- Wang etĀ al. (2017) Wang, X., Jones, T.Ā A., Treu, T., etĀ al. 2017, ApJ, 837, 89
- Wuyts etĀ al. (2016) Wuyts, E., Wisnioski, E., Fossati, M., etĀ al. 2016, ApJ, 827, 74
- Yuan etĀ al. (2011) Yuan, T.Ā T., Kewley, L.Ā J., Swinbank, A.Ā M., Richard, J., & Livermore, R.Ā C. 2011, ApJ, 732, L14
- Zaritsky etĀ al. (1994) Zaritsky, D., Kennicutt, RobertĀ C., J., & Huchra, J.Ā P. 1994, ApJ, 420, 87
- Zheng etĀ al. (2014) Zheng, W., Shu, X., Moustakas, J., etĀ al. 2014, ApJ, 795, 93
Appendix A Additional figures for A2744-YD4
Appendix B Additional figures for BDF-3299
Appendix C Additional figures for COSMOS24108